OpenStudy (anonymous):

find the common ratio for the following geometric sequence.... 2,1,0.5,0.25..

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Cr = 0.5

OpenStudy (anonymous):

1/2=.5/1=1/2 is the common ratio.